# Surrogate Modeling of Trajectory Map-matching in Urban Road Networks using Transformer Sequence-to-Sequence Model

### v2

@@ Abstract
Large-scale geolocation telematics data acquired from connected vehicles has the potential to significantly enhance mobility infrastructures and operational systems within smart cities. To effectively utilize this data, it is essential to accurately match the geolocation data to the road segments. However, this matching is often not trivial due to the low sampling rate and errors exacerbated by multipath effects in urban environments. Traditionally, statistical modeling techniques such as Hidden-Markov models incorporating domain knowledge into the matching process have been extensively used for map-matching tasks. However, rule-based map-matching tasks are noise-sensitive and inefficient in processing large-scale trajectory data. Deep learning techniques directly learn the relationship between observed data and road networks from the data, often without the need for hand-crafted rules or domain knowledge. This renders them an efficient approach for map-matching large-scale datasets and more robust to the noise. This paper introduces a deep-learning model, specifically the transformer-based encoder-decoder model, to perform as a surrogate for offline map-matching algorithms. The encoder-decoder architecture initially encodes the series of noisy GPS points into a representation that automatically captures autoregressive behavior and spatial correlations between GPS points. Subsequently, the decoder associates data points with the road network features and thus transforms these representations into a sequence of road segments. The model is trained and evaluated using GPS traces collected in Manhattan, New York. Achieving an accuracy of 75%, transformer-based encoder-decoder models extensively employed in natural language processing presented a promising performance for translating noisy GPS data to the navigated routes in urban road networks.
@@
map-matching, telematics data, GPS, GNSS, the transformer, sequence-to-sequence modeling, encoder-decoder, deep Learning.
@@ I INTRODUCTION
The emergence of smart cities and Intelligent Transportation Systems (ITS) marks a significant advancement in urban mobility planning and management [1]. Leveraging advanced technologies, these systems propel the development of transportation infrastructures [2], enhance traffic safety [3, 4, 5], contribute to the growth of intelligent and sustainable cities [6]. Central to this paradigm are Connected Vehicles (CVs), which enable Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) communication, creating a network for intelligent urban navigation [7]. The ubiquity of communication technologies has facilitated the collection of large amounts of real-world data, ranging from vehicle speed and location to traffic conditions. Leveraging the knowledge acquired from connected vehicles opens opportunities to gain new insights into driving behavior that enhances traffic predictions [8] and traffic safety [5]. Among the pool of data gathered by connected vehicles, the trajectory data derived from geolocation information holds numerous applications encompassing monitoring transportation systems and urban infrastructure, safety management [9, 10, 5], public transit efficiency improvement [11, 12], ride-hailing services [13], travel time prediction for operational systems [14, 15], and optimizing emergency response systems [16, 17]. By tracking the movement of vehicles, trajectory data facilitates route optimization and the identification of congestion zones and enhances the efficiency of urban road networks. Markovi et al. review trajectory data applications in different domains such as modeling human behavior, designing public transit, traffic performance measurement and prediction, environment, and safety [18].
Nevertheless, accurately deducing vehicle trajectories from noisy geolocation data obtained through Global Navigation Satellite Systems (GNSS) has become a key concern. In dense urban environments such as New York City, the abundance of tall buildings and narrow streets introduces complexities that challenge GNSS data collection. Obstacles like signal loss, reflections, and multipath interference introduce a significant source of error in GPS data, compromising the reliable and accurate representation of vehicular trajectory [19]. Figure 1 illustrates a multipath signal and compares the resultant erroneous path against the actual route on the map.
@@
Moreover, the quality of recovered routes is compromised by low sampling rates, measurement errors, and atmospheric conditions. Therefore, one of the major challenges in GPS trajectory data utilization pertains to handling noisy and imprecise GPS records.
Map-matching is a crucial process in transportation and geospatial data analysis that assigns measured GPS locations to specific road segments or locations on a map. Map-matching algorithms mostly rely on a combination of sensor data, including GPS coordinates, speed, and heading, and map data, such as road network topology and lane configurations. These algorithms employ a wide range of techniques spanning from statistical techniques to data-driven approaches, encompassing geometric algorithms, probabilistic models, and various advanced models [20, 21]. These techniques are employed to estimate the accurate positions and, consequently, to identify the most probable path or route taken by a vehicle or individual.
@@
Refining noisy GPS points to align them with consecutive road segments through advanced map-matching algorithms primarily involves a series of sophisticated and often challenging rules and techniques. An essential preliminary step, shared by many algorithms, involves the selection of candidate road segments for each observed GPS point (Figure 1 shows an example), then one seeks the most likely path among the set of all possible routes. Typically, road segments that fall within a confidence region often represented as an ellipse or a circle centered around the GPS point, are chosen as candidate roads. The dimension of the confidence region is determined based on the positioning error statistics, which is either a fixed distance threshold or an adaptive region [22]. How the path is estimated from the candidate segments varies depending on the particular map-matching algorithm being used. Different algorithms employ their unique strategies for choosing the most likely path based on the set of candidate segments.
However, with the increasing number of vehicles equipped with GPS devices that collect massive data, there is a growing need for a more efficient and scalable map-matching solution. Traditional map-matching tasks often involve a series of complex rules and data processing, making them inefficient, especially when confronted with large datasets or intricate algorithms. Deep learning surrogate models offer a promising solution to this challenge. In the context of map matching, a deep learning surrogate model refers to a neural network model crafted to replicate complex and often computationally demanding map-matching algorithms. Its purpose is to serve as a substitute for the original matching procedure, primarily aiming to accelerate and streamline the task while enhancing accuracy. By harnessing the wealth of historical trajectory data, data-driven algorithms are capable of discerning underlying mobility patterns and noise structures reflected by the data. Therefore, deep learning models autonomously discover the mapping function between input data (e.g., GPS coordinates) and output data (e.g., matched road segments) in an end-to-end manner. This task simplification reduces the necessity for extensive feature engineering, facilitating the alignment of GPS points with the road segments. Therefore, it yields a reliable tool for identifying trajectories within the road network.
This paper introduces a deep-learning model, in particular the transformer-based encoder-decoder model, to serve as a surrogate for advanced, rule-based map-matching algorithms. In particular, the transformer model that utilizes a self-attention mechanism in an encoder-decoder architecture [23] is trained to map the input sequence of GPS points onto a sequence of road segments. In the encoder-decoder architecture, the encoder initially encodes the noisy GPS points into a representation that captures the autoregressive behavior and spatial correlation between GPS points. Subsequently, the decoder captures the relationship between these GPS points and the road network features and thus transforms this representation into a sequence of road segments. The proposed deep-learning sequence-to-sequence model is a powerful technique capable of learning the most complex spatial-temporal characteristics of the data and the noise pattern across the road network from a large amount of trajectory data.
@@ III-A Map-matching preliminaries
This section outlines the definitions of the key components of the map-matching task and describes the format of the raw and matched data for the purpose of this paper.
Map-matching is a technique designed to determine the sequence of connected road segments on a digital road network map that best approximates the actual route taken by a vehicle, even in the presence of errors and noise in the observed GPS data. This task is challenging, especially in an urban environment, due to the significant positioning errors and the complex, dense road map. Map-matching task can be categorized into two types: online and offline matching. Online map-matching algorithms handle streaming GPS data in real time, providing immediate feedback on the moving vehicle’s current position relative to the map. In contrast, offline map-matching algorithms process historical trajectory data after it has been fully collected, allowing for more accurate and robust matching decisions by utilizing the entire dataset. This study proposes an offline map-matching algorithm that takes a complete sequence of observed coordinates in raw GPS data as input and produces a segment-based trajectory as the output. However, the low sampling rate in data collection leads to a sparse input representation. Consequently, no direct one-to-one correspondence exists between observed GPS points and inferred road segments.
@@
Raw GPS Trajectory is a collection of timestamped, noisy data points recorded by a GPS device at a frequency determined by the device setting. The sampling rate may vary depending on the purpose of the data collection and the capacity of the data storage infrastructure, resulting in either low or high sampling trajectory data. The collected data points typically include information about the vehicle’s position, such as latitude, longitude, altitude, and vehicle speed, as well as additional information, such as the number of satellites with which the GPS device communicates during location sampling. For the map-matching task in this study, the raw GPS trajectory is defined as a sequence order of latitude and longitude pairs, denoted by $T=\left\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})\right\}$. Here, $(x_{i},y_{i})$ represents the coordinates of the location points, and $n$ is the number of collected points along the route.
Segment-based Trajectory is a chronological sequence of connected road segments that accurately represents the vehicle’s path. In map-matching, a “road segment” is a portion of a digital road map corresponding to a particular stretch of road in the actual road network. When a GPS device records a vehicle’s position, the recorded latitude and longitude points are not perfectly aligned with the road segments on the digital map due to errors and noise.
Digital Map. A digital map or digital road network is a form of map data that is typically represented by a directed graph. The set of nodes represents the intersections or endpoints of the road segments, while the set of directed edges connects these nodes. In this graph representation, each directed edge corresponds to a road segment, with its direction indicating the allowed direction of traffic along the segment. By using this graph representation, map-matching algorithms accurately identify the segment-based trajectory from raw GPS trajectory.
@@ IV-A Problem Statement
This paper aims to develop a surrogate model for the map-matching tasks. The goal is to create a deep learning model that estimates the actual route taken by a vehicle in the form of a sequence of road segments when provided with a sequence of sparse, raw GPS coordinates. A seq2seq model is a well-suited approach to achieve this objective. This study employs a transformer-based encoder-decoder model to predict the most likely route based on a sequence of spatially unique grid cells onto which the noisy raw coordinates are projected.

### v4

# NLP-enabled Trajectory Map-matching in Urban Road Networks using a Transformer-based Encoder-decoder

@@ Abstract
Vehicular trajectory data, collected through geolocation telematics, is crucial for understanding mobility patterns in urban road networks. Map-matching, the process of aligning noisy and sparsely sampled GPS trajectories with digital road network map, is key to reconstructing accurate vehicle trajectories. Traditional methods rely on geometric proximity, topology, and shortest-path heuristics, overlooking two critical factors: (1) drivers often favor routes based on local road characteristics rather than shortest paths, revealing shared preferences that can be learned from large-scale trajectory data, and (2) GPS noise varies spatially mainly due to spatial variations in multipath effects. The interplay between these factors may make conventional methods less effective in complex real-world scenarios and may be more time-consuming for human experts to identify such heuristics and implement them. This study introduces an end-to-end and fully data-driven deep learning-based offline map-matching framework that formulates trajectory map-matching as a machine translation task, inspired by natural language processing (NLP). Specifically, it presents a transformer-based encoder-decoder map-matching framework that automatically learns contextual representations of noisy GPS points to infer drivers’ trajectory behavior and road network structures in an end-to-end manner. Trained on large-scale trajectory data, the proposed method improves path estimation accuracy in map-matching. Experiments on synthesized trajectories demonstrate that the proposed transformer-based map-matching framework outperforms conventional methods by integrating contextual awareness into trajectory analysis. Further evaluation on real-world GPS traces from Manhattan, New York, shows that our model achieves 75% accuracy in reconstructing navigated routes. These results demonstrate the advantage of transformer-based architectures in capturing real-world drivers’ trajectory behaviors, complex spatial dependencies and noise patterns, providing a more robust and scalable context-aware solution for map-matching. This study contributes to the broader vision of advancing trajectory-driven foundation models to enhance geospatial modeling and accelerate downstream urban mobility applications.
@@
Trajectory analysis, telematics data, GPS, GNSS, map-matching, transformer, sequential data modeling, encoder-decoder, deep Learning.
@@ I INTRODUCTION
The emergence of smart cities and Intelligent Transportation Systems (ITS) marks a significant advancement in urban mobility planning and management [1]. Leveraging advanced technologies, these systems propel the development of transportation infrastructures [2], improve traffic safety [3, 4, 5], and contribute to the growth of smart, sustainable cities [6]. Central to this paradigm are Connected Vehicles (CVs), which enable vehicle-to-everything (V2X) communication including Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) interactions, creating a network for intelligent urban navigation [7]. The ubiquity of the communication technologies has facilitated the collection of large amounts of real-world data, ranging from vehicle locations and speeds to traffic flow patterns. Analyzing this data offers insights into traffic patterns and drivers’ driving behavior, improving traffic modeling [8] and safety [5]. Among the pool of data gathered by connected vehicles, vehicular trajectory data holds numerous applications encompassing monitoring transportation systems and urban infrastructure, safety management [9, 10, 5], public transit efficiency improvement [11, 12], ride-hailing services [13], travel time prediction [14, 15], and optimizing emergency response systems [16, 17]. By tracking the movement of vehicles, trajectory data facilitates route optimization and the identification of congestion zones and enhances the efficiency of urban road networks. Markovi et al. review trajectory data applications in different domains such as modeling human behavior, designing public transit, traffic performance measurement and prediction, environment, and safety [18].
A vehicle trajectory represents the path a vehicle follows during a trip over time, typically expressed as a sequence of timestamped geolocation points. These points are collected by GPS receivers installed in the vehicle, which continuously record its position at specific time intervals. These receivers determine the vehicle’s position by communicating with a network of satellites orbiting Earth. However, in urban environments, the presence of tall buildings creates an “urban canyon” effect, which introduces errors in position calculations and consequently degrades the accuracy of GPS points. A major factor contributing to this issue is the multipath effect, where GPS signals reflect off surfaces such as buildings and reach the receiver via indirect paths. This causes incorrect position calculations due to the additional travel time of the reflected signals. Figure 1 shows an schematic example of the multipath effect and compares the resulting erroneous trajectory with the actual route on a map.
@@
In addition to the multipath effect, signal blockage and limited satellite visibility further degrade the accuracy of GPS trajectory data in urban settings [19]. Moreover, the dense and narrow street layouts of urban areas make it even more challenging to reconstruct the true path from a sequence of noisy GPS points. This challenge is further compounded when the GPS sampling rate is low, leading to a trajectory represented by a sparse sequence of noisy data points.
Therefore, a key challenge in utilizing trajectory data for downstream applications is the preprocessing of noisy and sparse GPS records. Map-matching techniques are used to align raw GPS points with the underlying road network, correcting positioning errors and inferring road segments between consecutive GPS points to eventually reconstruct the most likely vehicle trajectory. These techniques utilize a broad range of algorithms, from statistical to data-driven approaches, including geometric methods, probabilistic models, and advanced machine learning models [20, 21].
@@
A preliminary step in the majority of advanced map-matching algorithms involves the selection of candidate road segments for each observed GPS point (Figure 1 shows an example). Typically, road segments that fall within a search region, often represented as an ellipse or a circle centered around the GPS point, are chosen as candidate roads. The dimension of the search region is determined based on the positioning error statistics, which is either a fixed distance threshold or an adaptive region [22]. The algorithm then determines the most likely path by iteratively evaluating the set of all possible routes formed from the candidate segments. Path estimation from candidate road segments, i.e., identifying the most likely route, varies across map-matching algorithms, which apply geometric, topological, and probabilistic criteria. For instance, probabilistic models, such as Hidden Markov Models (HMMs), aim to maximize the joint emission and transition probabilities. Emission probability is the likelihood of observing noisy GPS points given that the vehicle is on a specific road segment. It is often calculated using the distance proximity and an estimate of the noise magnitude for the road network. Transition probability quantifies the conditional likelihood of transitioning from one road segment to the next. It is usually determined based on the difference between the observed, potentially erroneous trajectory and the shortest path between consecutive candidate points [23].
Path estimation in conventional map-matching algorithms primarily relies on geometric proximity, topology, and shortest-path heuristics, overlooking two critical factors that makes them less effective for map-matching sparse noisy data: (1) drivers often favor routes based on local road characteristics rather than shortest paths, revealing shared preferences that can be learned from large-scale trajectory data, and (2) GPS noise varies spatially mainly due to spatial variations in multipath effects. Street layouts and surrounding buildings contribute to the multipath effect, causing noise levels to fluctuate based on urban density. As illustrated in Figure 9, noise has a spatially non-uniform pattern which is high in dense urban areas with tall buildings and narrow streets. Most probabilistic map-matching approaches assume uniform GPS noise, making them noise sensitive. The interplay between drivers’ trajectory behavior, road network structure, and spatially varying noise makes conventional methods less effective in the real-world where data is typically noisy and sparsely collected. Therefore, in path estimation in the trajectory map-matching, context-aware approaches [24], are essential for effectively managing varying noise and drivers’ trajectory patterns.
Recent advancements in data-driven approaches, particularly deep learning and large-scale neural network models, have demonstrated remarkable success in learning contextual structures and capturing complex patterns across diverse domains, from drug discovery [25] to computer vision and natural language processing by leveraging large domain-specific datasets. Similarly, in geospatial modeling, these models effectively capture intricate mobility patterns by learning relationships between road network structure, the built environment as a key driver of spatial noise variation, and drivers’ navigation preferences across different locations. By autonomously learning structural features of the road network and uncovering hidden contextual factors influencing route choices and noise patterns, these models integrate this knowledge in an end-to-end manner to estimate the vehicular trajectory.
In this study, we draw an analogy between trajectory map-matching and the translation task in natural language processing (NLP). We conceptualize vehicle trajectories within large spatial datasets as analogous to sentences within a text corpus. Just as NLP models learn a variety of contextual patterns from large-scale language data to perform downstream tasks such as translation, similar architectures can be leveraged to extract contextual factors relevant to vehicle trajectories from large trajectory datasets. Therefore, we frame trajectory map-matching as a machine translation problem. As Figure 2 shows, in the former, a sequence of noisy, sparse GPS points is translated to an ordered sequence of road segments, while in the latter, a sequence of words in one language is translated to a sequence of words in another language. A similar analogy is presented by [26], defining map-matching as trajectory imputation by identifying its missing points, similar to filling in missing words in a sentence. Their work highlights the effectiveness of NLP models, particularly Bidirectional Encoder Representations from Transformers (BERT) [27], for the trajectory map-matching task.
In a similar fashion, this paper employs an encoder-decoder architecture with a transformer that leverages attention mechanisms [28] for the map-matching task. Trained on a large trajectory dataset, the encoder learns dense vector embeddings of tokenized, discretized GPS points (Figure 3 presents an example of discretized GPS points), capturing contextual structures such as spatial correlations. Meanwhile, the decoder learns embeddings of tokenized road segments, modeling the relationships between GPS points and road network features. Once trained, the model infers the sequence of road segments that best represents the traveled path given noisy, sparse GPS points. The rationale behind using the proposed deep learning sequence-to-sequence model is their capability to capture complex spatiotemporal dependencies, contextual factors influencing drivers’ trajectory behavior, and spatial noise variations across the road network by leveraging large-scale trajectory data.
Unlike traditional approaches such as Hidden Markov Models (HMMs) and other rule-based methods, transformers leverage embedding representations and attention mechanisms to learn relationships within and between GPS points and road network structures. This allows them to capture both local road network features and global contextual dependencies without relying on predefined transition probabilities or manually crafted rules. Consequently, transformers can adapt to varying noise levels across regions and account for drivers’ trajectory behaviors, enabling them to integrate information effectively when inferring missing road segments and accurately reconstructing the navigation path from sparse, noisy location points.
Moreover, large-scale neural network models, such as transformers, have enabled the development of pretrained foundation models across various domains, from natural language processing [29] to medicine [30]. When trained on extensive trajectory datasets from cities with diverse road structures, mobility patterns, and GPS noise distributions, these models have the potential to evolve into scalable, region-agnostic tools for trajectory analysis. However, realizing this adaptability requires further investigation. Advancing in this direction aligns with the vision of a universal trajectory foundation model, which could automate urban mobility analytics and accelerate trajectory analysis for a broad range of downstream mobility and geospatial applications.
Figure 2: NLP-enabled trajectory map-matching: recovering a complete, ordered sequence of connected road segments from a noisy, sparse, and discretized sequence of GPS points.
@@ III Map-matching preliminaries
This section outlines the definitions of the key components of the map-matching task and describes the data structure for the purpose of this paper.
Map-matching is a technique used to align noisy GPS coordinates with road segments from a digital road network map. Thereby, map-matching reconstructs the most likely path taken by a vehicle as a sequence of connected road segments. This task becomes particularly challenging in urban environments due to the complex and dense road layout, compounded by significant GPS positioning errors caused by signal obstructions and multipath effects, as well as the sparsity of GPS points in the trajectory. Map-matching tasks can be classified into two types: online and offline. Online map-matching algorithms process streaming GPS data in real time, providing immediate feedback on the moving vehicle’s current position relative to the map. In contrast, offline map-matching algorithms analyze historical trajectory data after it has been fully collected, focusing on improving path estimation accuracy by leveraging the entire observed, albeit erroneous, trajectory. The processed trajectory data is then used for various downstream applications.
Digital Map. A digital map or digital road network map is a structured representation of roadway data that stores both geographical and topological information, along with associated attributes. It is typically represented as a directed graph, denoted as $Graph=(V,R)$, where $V$ represents a set of nodes corresponding to the endpoints of road segments, and $R$ represents a set of directed edges connecting these nodes. Each directed edge, $r\in R$, corresponds to a road segment with its direction indicating the allowed direction of traffic along that segment. A road segment is a component of a digital road map corresponding to a particular stretch of road in the actual road network. Each edge typically includes additional attributes, such as road type, and speed limit. A digital map can be obtained directly as a graph from sources such as OpenStreetMap or constructed from publicly available regional spatial tabular data. For example, the geographic base information of New York City streets, stored in the LION shapefile maintained by the NYC Department of City Planning [58], can be used to generate a road network graph of the NYC.
@@
GPS-based Trajectory. Raw GPS-based trajectory is a collection of timestamped, noisy location points recorded by a GPS device at a frequency determined by the device setting. The sampling rate may vary depending on the purpose of the data collection and the capacity of the data storage infrastructure, resulting in either low or high sampling trajectory data. The collected data points typically include the vehicle’s position data, such as latitude, longitude, altitude, and speed, along with additional information, such as heading and the number of satellites the GPS device communicated with during location sampling. For the map-matching task in this study, the raw GPS-based trajectory is defined as a sequence order of latitude and longitude pairs, denoted by $T_{p}=\left\{p_{1}=(x_{1},y_{1}),p_{2}=(x_{2},y_{2}),\dots,p_{n}=(x_{n},y_{n})\right\}$. Here, $(x_{i},y_{i}),i=1,\dots,n$ represents the geographic coordinates of the location points, and $n$ is the number of collected points along the route.
Grid-based Trajectory. In spatial data analysis, the spatial region is often divided into a grid of discrete, non-overlapping cells, representing the trajectory as a sequence of visited cells rather than raw GPS coordinates. An example of a gridded region is shown in Figure 3. This discretization of GPS points is necessary for the modeling approach used in this paper. In this study, the entire region is partitioned into uniform square cells represented by the set $G$, where each cell has a unique identifier and a defined geometric shape. The total number of unique cells is denoted by $|G|$. Each GPS-based trajectory is transformed into a grid-based trajectory representation for further analysis. Let $f(.)$ be a function that maps the GPS point to the grid cell enclosing it based on the GPS point coordinates and the geometry of the grid cells. Using this function, the GPS-base trajectory of $T_{p}$ can be expressed as the ordered sequence of grid cells $T_{g}=\left\{g_{1},g_{2},\dots,g_{n}\right\}$, $g_{i}=f(p_{i})$, $g_{i}\in G$. Therefore, $g_{i}$ is the cell encompassing the GPS point $p_{i}$.
Segment-based Trajectory is a chronological sequence of connected road segments representing the vehicle’s path. Here, the segment-based trajectory is denoted by $T_{r}=\left\{r_{1},\dots,r_{j},\dots,r_{m}\right\}$ where $r_{j}\in R$ and $m$ is the number of connected road segments forming the segment-based trajectory.
Figure 3: A schematic representation of the input and output of the map-matching model.
Figure 4: Data structure: The left box shows a snippet of the raw trajectory consisting of $n$ noisy GPS points, denoted as ($p_{i}|i\in\left\{1,\dots,n\right\}$), sorted by timestamp. The middle box displays the corresponding grid cell IDs for each GPS point, denoted as ($g_{i}|i\in\left\{1,\dots,n\right\}$). The right box presents the map-matched trajectory, consisting of $m$ road segments, denoted as ($r_{j}|j\in\left\{1,\dots,m\right\}$).
This study introduces an offline map-matching framework that processes a raw, erroneous GPS-based trajectory, through first converting it to grid-based trajectory, and then outputs the most probable path as a segment-based trajectory. Figure 3 represents an example of the raw input (left), transformed input (middle), and the output (right). Figure 4 presents a snippet of the data structure for a sampled raw GPS trajectory, along with its grid-based and road segment-based representations. Since a low sampling rate in data collection results in a sparse GPS-based trajectory representation, the size of the segment-based trajectory may be equal or greater than that of the GPS-based trajectory, as multiple road segments can exist between two consecutive GPS points. For simplicity, in the following sections, $T_{g}$, $T_{r}$ are denoted as $g$ and $r$, respectively.
@@ IV-A Problem Statement
The goal of this paper is to create a deep learning model that estimates the actual route taken by a vehicle in the form of a sequence of road segments when provided with a sequence of sparse, raw GPS coordinates. A seq2seq model is a well-suited approach to achieve this objective. This study employs a transformer-based encoder-decoder model to predict the most likely route based on a sequence of spatially unique grid cells onto which the noisy raw coordinates are projected.